OpenStudy (kmullis6):

The table below represents the displacement of a turtle from its nest as a function of time:

OpenStudy (pawanyadav):

Y intercept is the value of function at x=0

OpenStudy (pawanyadav):

So can you tell what is y intercept here

OpenStudy (kmullis6):

@ganeshie8 Last question I have.. Thank you for all your help. Hope you don't mind one more. Part A: Find the slope. 510 - 285 = 225 6 - 3 = 3 225/3 = 75 y - 285 = 75(x - 3) y - 285 = 75x - 225 Add 285 to both sides y = 75x + 60 Part B: f(x) = 75x + 60 For Part C, I'm not sure how to do it.